Supposes a flexible cable is suspended between two towers that are 200 feet apart forms a curve whose equation is y=75(e^x/150 + e^-x/150). Calculate the length of the cable.
1
Expert's answer
2021-07-26T17:54:52-0400
"x_1=-100, x_2=100"

"y=75(e^{x\/150}+e^{-x\/150})"

"y'=\dfrac{1}{2}(e^{x/150}-e^{-x/150})"

"1+(y')^2=1+(\dfrac{1}{2}(e^{x/150}-e^{-x/150}))^2"

"=(\dfrac{1}{2}(e^{x/150}+e^{-x/150}))^2"

"L=\displaystyle\int_{-100}^{100}\sqrt{1+(y')^2}dx"

"=\displaystyle\int_{-100}^{100}\sqrt{(\dfrac{1}{2}(e^{x/150}+e^{-x/150}))^2}dx"


"=\displaystyle\int_{-100}^{100}\dfrac{1}{2}(e^{x/150}+e^{-x/150})dx"

"=2(\dfrac{1}{2})\displaystyle\int_{0}^{100}(e^{x/150}+e^{-x/150})dx"

"=150\big[e^{x/150}-e^{-x/150}\big]\begin{matrix}
 100 \\
 0
\end{matrix}"

"=150(e^{100\/150}-e^{-100\/150}-(1-1))"

"=150(e^{2/3}-e^{-2/3})\approx215.148\ m"